Dedicared Fun Corporation issued 2,000 shares of common stock and 400 shares of preferred stock for a lump sum of $72,000 cash.

  1. Give the entry for the issuance assuming the par value of the common stock was $5 and the fair value $30, and the par value of the preferred stock was $40 and the fair value $50. (Each valuation is on a per share basis and there are ready markets for each stock.)






  2. Give the entry for the issuance assuming the preferred stock has no ready market and the common stock has a fair value of $23 per share. Par values are the same as in part 1.
Answer #1
Fair value of Common Stock 60000 =2000*30
Fair value of Preferred Stock 20000 =400*50
Total fair value 80000
Value allocated to Common Stock 54000 =72000*60000/80000
Value allocated to Preferred Stock 18000 =72000*20000/80000
a
General Journal Debit Credit
Cash 72000
      Common Stock 10000 =2000*5
      Paid-in Capital in Excess of Par—Common Stock 44000 =54000-10000
      Preferred Stock 16000 =400*40
      Paid-in Capital in Excess of Par—Preferred Stock 2000 =18000-16000
b
General Journal Debit Credit
Cash 72000
      Common Stock 10000 =2000*5
      Paid-in Capital in Excess of Par—Common Stock 36000 =2000*(23-5)
      Preferred Stock 16000 =400*40
      Paid-in Capital in Excess of Par—Preferred Stock 10000